An open flask sitting in a lab refrigerator looks empty, but it is actually filled with a mixture of gases called air. If the flask volume is 1.00 L, and the air is at standard temperature and pressure, how many gaseous molecules does the flask contain?


1
Expert's answer
2021-11-10T00:29:44-0500

At standard temperature and pressure, any gas occupies "22.4l"

"\therefore" "1l" will be occupied by


"\frac{1}{22.4}\times1=0.04464\>moles"


But the number of particles in one mole of any substance is:

"6.023\times10^{23}"


Therefore, "0.04464 \,moles" will have


"\frac{0.04464}{1}\times(6.023\times10^{23}" ")"


"= 2.689\times10^{22}" Molecules